When three times a number is subtracted from 40, the result is 4 plus the number. What is the number?
Answer by anerson 25(3)

thanks this is my my possible solution to that equation:40-3n=n+4
-3n-n=4-40
-2n=-36
n=18
CHECK:40-3(18)=18+4
40-54=22